Consider the following information

Period 1 2 3 4 5

Forecast 63 72 86 71 60

Actual 57 60 70 75 70

The mean absolute deviation M AD is

1. 552

2. 9:6

3. 48

4. 98

5. 96


1
Expert's answer
2021-08-17T18:02:49-0400

"MAD = \frac{\sum^n_{i=1} |y_t -F_t|}{n} \\

MAD = \frac{|57-63|+|60-72|+|70-86|+|75-71|+|70-60|}{5} \\

= \frac{6+12+16+4+10}{5} \\

= 9.6"

Answer: option 2 is correct